OpenStudy (fatimab2000):

Will Medal !!!!!! Larry is using an online calculator to calculate the outputs f(n) for different inputs n. The ordered pairs below show Larry's inputs and the corresponding outputs displayed by the calculator. (1, 5), (2, 9), (3, 13), (4, 17) Which of the following functions best represents the rule that the calculator uses to display the outputs? f(n) = 5n - 1 f(n) = 5n + 1 f(n) = 4n + 1 f(n) = 4n - 1

OpenStudy (fatimab2000):

Help Please

OpenStudy (fatimab2000):

@jigglypuff314

OpenStudy (welshfella):

plug n = 1 into each function and see which one gives you 5

OpenStudy (welshfella):

so for the first function f(1) = 5(1) - 1 = 4 so what does that tell you?

OpenStudy (fatimab2000):

i just really need confirmation i think its C

OpenStudy (welshfella):

does see fit the ordered pair (1,5)?

OpenStudy (welshfella):

* C

OpenStudy (fatimab2000):

no

OpenStudy (fatimab2000):

yes

OpenStudy (anonymous):

Confirmation from Mathematica using a very small program. n ranges from 1 through 4. Table[ { n, 4n+1 }, { n, 4 } ] = { {1,5}, {2,9}, {3,13}, {4,17} }